### v1

@@
marginparsep has been altered.
topmargin has been altered.
marginparpush has been altered.
The page layout violates the ICML style.Please do not change the page layout, or include packages like geometry, savetrees, or fullpage, which change it for you. We’re not able to reliably undo arbitrary changes to the style. Please remove the offending package(s), or layout-changing commands and try again.
@@
Tong Wu $^{\dagger}$ Yang Liu*  Jun Bai*  Zixia Jia Shuyi Zhang  Ziyong Lin  Yanting Wang
@@ 1 Introduction
The advent of super-scale Large Language Models (LLMs), exemplified by Gemini 3 (gemini3), GPT-5 (gpt5), and DeepSeek-V3.2 (deepseek_v3.2), has shifted the frontier of AI from semantic fluency to deep, multi-step agentic reasoning. Despite the excitement of “deeper” test-time scaling that enables models to solve complex problems muennighoff2025s1simpletesttimescaling, the “wider” reasoning capacity to explore diverse trajectories in parallel emerges as the dominant requirement toward agentic AI (shen2025thinkingvsdoingagents, gemini2.5). The MapReduce paradigm has long underpinned distributed computing by separating task decomposition from trajectory aggregation (mapreduce, multiverse, wang2025llmtimesmapreducev2), yet its application to agentic language modeling remains a critical missing link in the evolution of open-source LLMs. Ideally, the model should internalize the collaborative breadth of multi-agent systems directly into an efficient, natively parallel architecture.
@@
Despite this clear imperative, existing implementations remain fragmented and present three critical deficiencies. First, Algorithmic and Architectural Incompatibility. Prevalent inference engines (sglang, vllm) and Reinforcement Learning (RL) algorithms (grpo) are ill-equipped for native branching: The former fails to control parallel branching and aggregation; the latter often clips the gradients of the special tokens that trigger those operations, preventing the model from learning strict structure. Second, Inefficient Hand-Crafted Parallelism. Although the intuitive advantage of parallel sampling lies in efficiency, early attempts on internalizing the parallelism parallel_r1, DBLP:journals/corr/abs-2509-04475/parathinker, zhao2025parallelsearch resort to hand-crafted divide-and-conquer rules via independent sampling. These methods fail to leverage shared Key-Value (KV) states, necessitating redundant recalculations for every branch and resulting in prohibitive linear latency costs ($O(N)$) that render the model impractical for real-time deployment. Third, Reliance on Supervised Distillation. Framework such as Multiverse (multiverse) successfully operationalizes native parallelism but depends heavily on supervised data distilled from stronger teacher models. While effective for compressing capabilities into smaller models, this dependence restricts the student to mimicking the teacher’s sequential reasoning topology force-fitted into a parallel format, imposing an "Intelligence Ceiling" that prevents it from novel, model-intrinsic parallel strategies necessary for super-intelligence.
@@
To address these challenges, we take the first step to explore the potential of LLMs to self-evolve parallel reasoning capabilities without reliance on external supervision and introduce Native Parallel Reasoner (NPR). Specifically, NPR employs a three-stage progressive training paradigm designed to transition the model from sequential emulation to genuine parallel cognition. In Stage 1 ($\S$2.2), we warm up a seed instruction-tuned model (e.g., Qwen3-4B-Instruct) to spontaneously discover valid parallel structures by applying standard DAPO (dapo) with a format-aware reward function, yielding a structured trajectory generator, NPR-Zero, which produces parallel-formatted outputs but still relies on sequential visibility, i.e., simulated parallelism. In Stage 2 ($\S$2.3), we bridge the gap to native parallel architecture by performing rejection sampling on NPR-Zero and conducting parallel warmup, which instills strict topological constraints using parallel positional encoding and attention mask as in multiverse, converting the sequential behavior of models into real parallel execution. Finally, Stage 3 ($\S$2.4) generalizes these capabilities beyond the initial distribution via Native-Parallel RL. This stage implements collision-free parallel rollouts using a novel Parallel-Aware Policy Optimization (PAPO) algorithm, which optimizes branching policies directly within the parallel execution graph, allowing the model to learn adaptive decomposition strategies through trial and error rather than imitation.
@@
  • Self-Distilled Data Efficacy: Our self-distilled datasets outperform previous teacher-generated trajectories in multiverse by an average of 10.1 points, validating the hypothesis to learn from native distributions.
@@ 1.1 Contributions
  • We re-engineered an NPR Engine, which redesigns core components of parallel execution, including radix-cache memory management, output-length control, and parallel-state branching based on Multiverse-Engine (multiverse). This eliminates the instability inherent in standard engines, providing a practical backend for native parallel RL training.
@@ 2.1 Preliminaries
To optimize the policy model within our reinforcement learning framework, we adopt objective functions based on DAPO (dapo). We first introduce the original DAPO update procedure. For each question-answer pair $(q,y)\sim\mathcal{D}$, the policy model $\pi_{\theta_{\text{old}}}$ first generates a group of responses $\{\hat{y}_{i}\}_{i=1}^{G}$. The objective function $\mathcal{J}(\theta)$ is then formulated as:
@@ 2.2 Stage 1: Format-follow Reinforcement Learning
To support adaptive decomposition and parallel reasoning during generation, we adopt a simplified “Map–Process–Reduce” schema inspired by Multiverse (multiverse) but with a leaner structure. Each parallel block begins with <guideline> ... </guideline>, which contains a set of <plan> ... </plan> entries that define the Map stage. The Process stage follows: each <step> ... </step> block executes one mapped subtask independently and in parallel. After all <step> blocks complete, a Reduce stage consolidates their outputs into a final summary wrapped by <takeaway> ... </takeaway>. This explicit tag-based format makes the decomposition, independent processing, and final aggregation easy to parse and verify in downstream training and evaluation.
While this schema provides a clear, learnable format for parallel reasoning, obtaining large-scale, high-quality training data for it remains challenging. Prior work such as Multiverse (multiverse) constructs large, multi-step synthetic pipelines and aggregates outputs from several state-of-the-art teacher models (e.g., Deepseek R1 (deepseek_r1) and Gemini 2.5 Pro (gemini2.0)) to overcome data scarcity. While effective, these multi-teacher pipelines add operational complexity, require access to strong external teachers, and incur substantial maintenance costs.
We adopt a simpler, self-improving approach. Starting from a single pretrained LLM, we apply DAPO (dapo) to induce the target native parallel-reasoning generation format without paired supervision or external teachers. Our reward function combines format and accuracy signals. For format: outputs that pass a format check receive a reward of 0.0; outputs that fail receive a penalty in (0.0, -2.0]. For accuracy: when the format check passes, correct answers yield +1.0 and incorrect answers yield -1.0. The checkpoint produced by this process (denoted NPR-Zero) is therefore optimized primarily to learn the required structured format; we then use its generations for large-scale self-distillation to build a synthetic corpus for downstream supervised fine-tuning (SFT).
@@ 2.3 Stage 2: Rejection Sampling and Parallel Warmup
$\begin{split}\mathcal{D}_{\text{accept}}=\{(q_{i},r^{i}_{j},a^{i}_{j})\mid i\leq N,j\leq K,\text{ s.t. }(r^{i}_{j},a^{i}_{j})\sim\pi_{\theta}(\cdot|q_{i}),\mathbb{1}_{\text{accept}}(r^{i}_{j},\hat{a}^{i}_{j})=1\}.\end{split}$ (4)
@@
To support structured parallel generation, we adopt the core design of Multiverse Attention (multiverse) when constructing both the parallel attention mask and the corresponding positional encoding (Algorithm˜1 and Algorithm˜2). This design enables multiple reasoning paths to coexist within a single forward pass while allowing fast adaptation from only a few examples. It also permits efficient KV-cache reuse for the shared context inside the NPR Engine (§2.5), reducing inference overhead. Furthermore, to ensure the model can emit the required structural tags, we initialize a set of special tokens that correspond to these tags and expose them during the cold-start training stage.
@@ 2.4 Stage 3: Native-parallel RL
Figure 3: Comparison of GRPO-style RL (grpo) and Parallel-Aware Policy Optimization.
Below we summarize the practical modifications we make to standard RL (dapo) to respect parallel semantics and stabilize training.
@@
Existing inference engines (vllm, sglang) do not enforce strict parallel semantics, so they can produce malformed trajectories. We therefore sample rollouts using our NPR-Engine (§2.5), which guarantees that every generated trajectory follows the intended Map–Process–Reduce flow.
@@
Because format-violating samples are removed before optimization, group-level variance collapses, which makes relative (group) advantages ineffective. We adopt a Lite-PPO (lite_ppo) style advantage but replace group-level variance with batch-level variance. For each sample $i$ and token $t$ we compute
@@
Special tokens [^4] The tags that control parallel branching and merging. are critical to maintain parallel semantics. Token-level clipping that suppresses gradients for these tokens breaks the learned structure, so we remove clip-masking and ensure special tokens always receive gradients.However, removing clip-masking makes importance-sampling ratios in PPO (ppo) unstable. To avoid unstable reweighting, we eliminate importance sampling and adopt a strict on-policy objective. This both stabilizes training and speeds it up because we do not need to recompute historical log-probabilities.
@@ 2.5 Engineering Enhancement: NPR Engine
Multiverse’s parallel-generation engine (multiverse) supplies a powerful substrate for large-scale rollout based on SGLang [^5] https://github.com/sgl-project/sglang, but when exercised at production scale, it exposed a set of brittle implementation corners that undermine both correctness and RL stability. We implemented a compact set of engine-level mitigation to restore deterministic behavior, memory safety, and correct length accounting across high-throughput parallel rollouts, which together form the NPR-Engine used in our Parallel-RL pipeline.
@@ 3 Experiments
We build our experiments on the ORZ dataset (orz), which contains 57k problem–answer pairs. To ensure consistency across all stages of our pipeline, we sample a fixed subset of 8k examples from ORZ and use this for Stage 1 (§2.2), Stage 2 (§2.3), and Stage 3 (§2.4).
@@
Our models are based on Qwen3-4B-Instruct-2507 and Qwen3-4B (non-thinking mode) (DBLP:journals/corr/abs-2505-09388/qwen3). We intentionally avoid the thinking-mode variant because it cannot be trained with standard supervised fine-tuning (§4.6).
@@ 3.2 Evaluation Setup
We assess the effectiveness and generalization ability of NPR across a diverse suite of reasoning benchmarks. For relatively small-scale datasets such as AIME24 (maa_aime2024), AIME25 (maa_aime2025), HMMT25 (balunovic2025matharena), and AMC23 (maa_amc12_problems_solutions), we report avg@8, which better reflects performance when multiple sampled solutions are available. For larger or more heterogeneous benchmarks including OlympiadBench (he-etal-2024-olympiadbench), Minerva-Math (lewkowycz2022solving), ZebraLogic (zebralogic), and MATH500 (math500), we follow the standard single-answer setting and report avg@1.
@@
  • Open Sequential Reasoners: Qwen2.5-32B-Instruct (yang2024qwen25), Qwen3-4B (DBLP:journals/corr/abs-2505-09388/qwen3) (without thinking mode), and Qwen3-4B-Instruct-2507.
  • Recent Parallel Reasoners: Multiverse (multiverse) models, including Multiverse-32B and our reproduced Multiverse-4B built on Qwen3-4B-Instruct-2507.
@@ 5 Related Work
Parallel reasoning improves reasoning efficiency and robustness by exploring multiple reasoning paths simultaneously, unlike standard sequential reasoning which is prone to early commitment errors (the “prefix trap”) and lacks self-correction, leading to suboptimal solutions and slow inference due to its strictly step-by-step generation process (DBLP:journals/corr/abs-2510-12164/parallelreasoning). Early methods, such as Best-of-N (DBLP:journals/corr/abs-2110-14168/bon) and Self-Consistency (DBLP:conf/iclr/0002WSLCNCZ23/self-consistency), select the most scored or consistent output from independent paths but are not end-to-end optimized. Search-based approaches like Tree-of-Thought (Tot), Graph-of-Thought (DBLP:conf/aaai/BestaBKGPGGLNNH24/graph-of-thoughts), Monte Carlo Tree Search (DBLP:journals/corr/abs-2405-00451/mcts) further explore reasoning trees but rely on hand-designed structures and external verifiers, limiting flexibility and scalability. To further improve the adaptability and flexibility of parallel reasoning operations, recent work strives through learning approaches. One line of work adopts the SFT paradigm—for example, Multiverse (multiverse), ParaThinker (DBLP:journals/corr/abs-2509-04475/parathinker), and SSFT (DBLP:journals/corr/abs-2510-05132/globeforking)—which guide model learning through parallel reasoning paths derived from the sequential trajectories of more powerful large reasoning models (LRMs). However, such pure imitation limits the model’s ability to discover novel reasoning patterns. Another line of work enhances parallel reasoning capabilities through reinforcement learning (RL), such as APR (DBLP:journals/corr/abs-2504-15466/APR) and Parallel-R1 (parallel_r1). However, these methods either demonstrate effectiveness only on toy tasks or still depend on supervised data from other reasoning models to bootstrap the RL process.
@@
Reinforcement learning (RL) has become an important tool for enhancing the reasoning capabilities of large language models (LLMs) in recent years (DBLP:journals/corr/abs-2403-04642, DBLP:journals/corr/abs-2509-08827, tokenswift, seek, rulereasoner, routerlens, cream). Early and widely adopted approaches—such as RL from human feedback (RLHF)—optimize outcome-level rewards derived from human preferences or task-level correctness (simpo). These methods improve alignment and robustness in general generation tasks but provide only coarse control over intermediate reasoning trajectories. Then, research shifts toward process-aware RL, where step-level reward modeling offer denser and more interpretable supervision (oai_prm, DBLP:conf/acl/ZhangZWZLYLZL25, DBLP:journals/corr/abs-2504-16828). Those process-level feedback, however, suffer from subjectivity, high annotation cost, and unstable optimization due to ambiguous or unverifiable intermediate signals. A further evolution leads to Reinforcement Learning with Verifiable Reward (RLVR), which replaces opaque reward models with explicit, auditable verifiers (e.g., logical checkers, rule-based graders, or formal validators) (grpo, DBLP:journals/corr/abs-2502-14768, dapo, gspo). Compared with conventional reward modeling, RLVR provides objectivity, reproducibility, and stronger correctness guarantees, making it particularly suited for reasoning tasks where outputs are verifiable (e.g., math, programming, or factual QA). Moreover, RLVR reduces human labeling costs by leveraging deterministic verifiers as reward oracles.

### v2

@@
[…]
@@ 1 Introduction
The advent of super-scale Large Language Models (LLMs), exemplified by Gemini 3 (Pichai et al., 2025), GPT-5 (OpenAI, 2025), and DeepSeek-V3.2 (DeepSeek, Inc., 2025), has shifted the frontier of AI from semantic fluency to deep, multi-step agentic reasoning. Despite the excitement of “deeper” test-time scaling that enables models to solve complex problems Muennighoff et al. (2025), the “wider” reasoning capacity to explore diverse trajectories in parallel emerges as the dominant requirement toward agentic AI (Shen et al., 2025, Comanici et al., 2025). The MapReduce paradigm has long underpinned distributed computing by separating task decomposition from trajectory aggregation (Dean and Ghemawat, 2008, Yang et al., 2025a, Wang et al., 2025a), yet its application to agentic language modeling remains a critical missing link in the evolution of open-source LLMs. Ideally, the model should internalize the collaborative breadth of multi-agent systems directly into an efficient, natively parallel architecture.
@@
Despite this clear imperative, existing implementations remain fragmented and present three critical deficiencies. First, Algorithmic and Architectural Incompatibility. Prevalent inference engines (Zheng et al., 2024, Kwon et al., 2023) and Reinforcement Learning (RL) algorithms (Shao et al., 2024) are ill-equipped for native branching: The former fails to control parallel branching and aggregation; the latter often clips the gradients of the special tokens that trigger those operations, preventing the model from learning strict structure. Second, Inefficient Hand-Crafted Parallelism. Although the intuitive advantage of parallel sampling lies in efficiency, early attempts on internalizing the parallelism Zheng et al. (2025a), Wen et al. (2025), Zhao et al. (2025) resort to hand-crafted divide-and-conquer rules via independent sampling. These methods fail to leverage shared Key-Value (KV) states, necessitating redundant recalculations for every branch and resulting in prohibitive linear latency costs ($O(N)$) that render the model impractical for real-time deployment. Third, Reliance on Supervised Distillation. Framework such as Multiverse (Yang et al., 2025a) successfully operationalizes native parallelism but depends heavily on supervised data distilled from stronger teacher models. While effective for compressing capabilities into smaller models, this dependence restricts the student to mimicking the teacher’s sequential reasoning topology force-fitted into a parallel format, imposing an "Intelligence Ceiling" that prevents it from novel, model-intrinsic parallel strategies necessary for super-intelligence.
@@
To address these challenges, we take the first step to explore the potential of LLMs to self-evolve parallel reasoning capabilities without reliance on external supervision and introduce Native Parallel Reasoner (NPR). Specifically, NPR employs a three-stage progressive training paradigm designed to transition the model from sequential emulation to genuine parallel cognition. In Stage 1 ($\S$2.2), we warm up a seed instruction-tuned model (e.g., Qwen3-4B-Instruct) to spontaneously discover valid parallel structures by applying standard DAPO (Yu et al., 2025) with a format-aware reward function, yielding a structured trajectory generator, NPR-Zero, which produces parallel-formatted outputs but still relies on sequential visibility, i.e., simulated parallelism. In Stage 2 ($\S$2.3), we bridge the gap to native parallel architecture by performing rejection sampling on NPR-Zero and conducting parallel warmup, which instills strict topological constraints using parallel positional encoding and attention mask as in Yang et al. (2025a), converting the sequential behavior of models into real parallel execution. Finally, Stage 3 ($\S$2.4) generalizes these capabilities beyond the initial distribution via Native-Parallel RL. This stage implements collision-free parallel rollouts using a novel Parallel-Aware Policy Optimization (PAPO) algorithm, which optimizes branching policies directly within the parallel execution graph, allowing the model to learn adaptive decomposition strategies through trial and error rather than imitation.
@@
  • Self-Distilled Data Efficacy: Our self-distilled datasets outperform previous teacher-generated trajectories in Yang et al. (2025a) by an average of 10.1 points, validating the hypothesis to learn from native distributions.
@@ 1.1 Contributions
  • We re-engineered an NPR Engine, which redesigns core components of parallel execution, including radix-cache memory management, output-length control, and parallel-state branching based on Multiverse-Engine (Yang et al., 2025a). This eliminates the instability inherent in standard engines, providing a practical backend for native parallel RL training.
@@ 2.1 Preliminaries
To optimize the policy model within our reinforcement learning framework, we adopt objective functions based on DAPO (Yu et al., 2025). We first introduce the original DAPO update procedure. For each question-answer pair $(q,y)\sim\mathcal{D}$, the policy model $\pi_{\theta_{\text{old}}}$ first generates a group of responses $\{\hat{y}_{i}\}_{i=1}^{G}$. The objective function $\mathcal{J}(\theta)$ is then formulated as:
@@ 2.2 Stage 1: Format-follow Reinforcement Learning
To support adaptive decomposition and parallel reasoning during generation, we adopt a simplified “Map–Process–Reduce” schema inspired by Multiverse (Yang et al., 2025a) but with a leaner structure. Each parallel block begins with <guideline> ... </guideline>, which contains a set of <plan> ... </plan> entries that define the Map stage. The Process stage follows: each <step> ... </step> block executes one mapped subtask independently and in parallel. After all <step> blocks complete, a Reduce stage consolidates their outputs into a final summary wrapped by <takeaway> ... </takeaway>. This explicit tag-based format makes the decomposition, independent processing, and final aggregation easy to parse and verify in downstream training and evaluation.
While this schema provides a clear, learnable format for parallel reasoning, obtaining large-scale, high-quality training data for it remains challenging. Prior work such as Multiverse (Yang et al., 2025a) constructs large, multi-step synthetic pipelines and aggregates outputs from several state-of-the-art teacher models (e.g., Deepseek R1 (Guo et al., 2025) and Gemini 2.5 Pro (Google, 2025)) to overcome data scarcity. While effective, these multi-teacher pipelines add operational complexity, require access to strong external teachers, and incur substantial maintenance costs.
We adopt a simpler, self-improving approach. Starting from a single pretrained LLM, we apply DAPO (Yu et al., 2025) to induce the target native parallel-reasoning generation format without paired supervision or external teachers. Our reward function combines format and accuracy signals. For format: outputs that pass a format check receive a reward of 0.0; outputs that fail receive a penalty in (0.0, -2.0]. For accuracy: when the format check passes, correct answers yield +1.0 and incorrect answers yield -1.0. The checkpoint produced by this process (denoted NPR-Zero) is therefore optimized primarily to learn the required structured format; we then use its generations for large-scale self-distillation to build a synthetic corpus for downstream supervised fine-tuning (SFT).
@@ 2.3 Stage 2: Rejection Sampling and Parallel Warmup
$\begin{split}\mathcal{D}_{\text{accept}}=\{(q_{i},r^{i}_{j},\hat{a}^{i}_{j})\mid i\leq N,j\leq K,\text{ s.t. }(r^{i}_{j},\hat{a}^{i}_{j})\sim\pi_{\theta}(\cdot|q_{i}),\mathbb{1}_{\text{accept}}(r^{i}_{j},\hat{a}^{i}_{j})=1\}.\end{split}$ (4)
@@
To support structured parallel generation, we adopt the core design of Multiverse Attention (Yang et al., 2025a) when constructing both the parallel attention mask and the corresponding positional encoding (Algorithm˜1 and Algorithm˜2). This design enables multiple reasoning paths to coexist within a single forward pass while allowing fast adaptation from only a few examples. It also permits efficient KV-cache reuse for the shared context inside the NPR Engine (§2.5), reducing inference overhead. Furthermore, to ensure the model can emit the required structural tags, we initialize a set of special tokens that correspond to these tags and expose them during the cold-start training stage.
@@ 2.4 Stage 3: Native-parallel RL
Figure 3: Comparison of GRPO-style RL (Shao et al., 2024) and Parallel-Aware Policy Optimization.
Below we summarize the practical modifications we make to standard RL (Yu et al., 2025) to respect parallel semantics and stabilize training.
@@
Existing inference engines (Kwon et al., 2023, Zheng et al., 2024) do not enforce strict parallel semantics, so they can produce malformed trajectories. We therefore sample rollouts using our NPR-Engine (§2.5), which guarantees that every generated trajectory follows the intended Map–Process–Reduce flow.
@@
Because format-violating samples are removed before optimization, group-level variance collapses, which makes relative (group) advantages ineffective. We adopt a Lite-PPO (Liu et al., 2025a) style advantage but replace group-level variance with batch-level variance. For each sample $i$ and token $t$ we compute
@@
Special tokens [^4] The tags that control parallel branching and merging. are critical to maintain parallel semantics. Token-level clipping that suppresses gradients for these tokens breaks the learned structure, so we remove clip-masking and ensure special tokens always receive gradients.However, removing clip-masking makes importance-sampling ratios in PPO (Schulman et al., 2017) unstable. To avoid unstable reweighting, we eliminate importance sampling and adopt a strict on-policy objective. This both stabilizes training and speeds it up because we do not need to recompute historical log-probabilities.
@@ 2.5 Engineering Enhancement: NPR Engine
Multiverse’s parallel-generation engine (Yang et al., 2025a) supplies a powerful substrate for large-scale rollout based on SGLang [^5] https://github.com/sgl-project/sglang, but when exercised at production scale, it exposed a set of brittle implementation corners that undermine both correctness and RL stability. We implemented a compact set of engine-level mitigation to restore deterministic behavior, memory safety, and correct length accounting across high-throughput parallel rollouts, which together form the NPR-Engine used in our Parallel-RL pipeline.
@@ 3 Experiments
We build our experiments on the ORZ dataset (Hu et al., 2025), which contains 57k problem–answer pairs. To ensure consistency across all stages of our pipeline, we sample a fixed subset of 8k examples from ORZ and use this for Stage 1 (§2.2), Stage 2 (§2.3), and Stage 3 (§2.4).
@@
Our models are based on Qwen3-4B-Instruct-2507 and Qwen3-4B (non-thinking mode) (Yang et al., 2025b). We intentionally avoid the thinking-mode variant because it cannot be trained with standard supervised fine-tuning (§4.6).
@@ 3.2 Evaluation Setup
We assess the effectiveness and generalization ability of NPR across a diverse suite of reasoning benchmarks. For relatively small-scale datasets such as AIME24 (Mathematical Association of America, 2024), AIME25 (Mathematical Association of America, 2025), HMMT25 (Balunović et al., 2025), and AMC23 (Mathematical Association of America, 2023), we report avg@8, which better reflects performance when multiple sampled solutions are available. For larger or more heterogeneous benchmarks including OlympiadBench (He et al., 2024), Minerva-Math (Lewkowycz et al., 2022), ZebraLogic (Lin et al., 2025), and MATH500 (Hendrycks et al., 2021), we follow the standard single-answer setting and report avg@1.
@@
  • Open Sequential Reasoners: Qwen2.5-32B-Instruct (Qwen et al., 2024), Qwen3-4B (Yang et al., 2025b) (without thinking mode), and Qwen3-4B-Instruct-2507.
  • Recent Parallel Reasoners: Multiverse (Yang et al., 2025a) models, including Multiverse-32B and our reproduced Multiverse-4B built on Qwen3-4B-Instruct-2507.
@@ 5 Related Work
Parallel reasoning improves reasoning efficiency and robustness by exploring multiple reasoning paths simultaneously, unlike standard sequential reasoning which is prone to early commitment errors (the “prefix trap”) and lacks self-correction, leading to suboptimal solutions and slow inference due to its strictly step-by-step generation process (Wang et al., 2025b). Early methods, such as Best-of-N (Cobbe et al., 2021) and Self-Consistency (Wang et al., 2023), select the most scored or consistent output from independent paths but are not end-to-end optimized. Search-based approaches like Tree-of-Thought (Yao et al., 2023), Graph-of-Thought (Besta et al., 2024), Monte Carlo Tree Search (Xie et al., 2024) further explore reasoning trees but rely on hand-designed structures and external verifiers, limiting flexibility and scalability. To further improve the adaptability and flexibility of parallel reasoning operations, recent work strives through learning approaches. One line of work adopts the SFT paradigm—for example, Multiverse (Yang et al., 2025a), ParaThinker (Wen et al., 2025), and SSFT (Jia et al., 2025)—which guide model learning through parallel reasoning paths derived from the sequential trajectories of more powerful large reasoning models (LRMs). However, such pure imitation limits the model’s ability to discover novel reasoning patterns. Another line of work enhances parallel reasoning capabilities through reinforcement learning (RL), such as APR (Pan et al., 2025) and Parallel-R1 (Zheng et al., 2025a). However, these methods either demonstrate effectiveness only on toy tasks or still depend on supervised data from other reasoning models to bootstrap the RL process.
@@
Reinforcement learning (RL) has become an important tool for enhancing the reasoning capabilities of large language models (LLMs) in recent years (Havrilla et al., 2024, Zhang et al., 2025a, Wu et al., 2025, Li et al., 2025, Liu et al., 2025b, Bai et al., 2025, Wu et al., 2024). Early and widely adopted approaches—such as RL from human feedback (RLHF)—optimize outcome-level rewards derived from human preferences or task-level correctness (Meng et al., 2024). These methods improve alignment and robustness in general generation tasks but provide only coarse control over intermediate reasoning trajectories. Then, research shifts toward process-aware RL, where step-level reward modeling offer denser and more interpretable supervision (Lightman et al., 2024, Zhang et al., 2025b, Khalifa et al., 2025). Those process-level feedback, however, suffer from subjectivity, high annotation cost, and unstable optimization due to ambiguous or unverifiable intermediate signals. A further evolution leads to Reinforcement Learning with Verifiable Reward (RLVR), which replaces opaque reward models with explicit, auditable verifiers (e.g., logical checkers, rule-based graders, or formal validators) (Shao et al., 2024, Xie et al., 2025, Yu et al., 2025, Zheng et al., 2025b). Compared with conventional reward modeling, RLVR provides objectivity, reproducibility, and stronger correctness guarantees, making it particularly suited for reasoning tasks where outputs are verifiable (e.g., math, programming, or factual QA). Moreover, RLVR reduces human labeling costs by leveraging deterministic verifiers as reward oracles.